# Intumescent Polymer-Particle Hydrogels Expand the Retardancy Window for Wildfire Defense

### v1

@@ Abstract
The 2025 Los Angeles wildfires caused widespread urban destruction and displacement, and severe economic losses, highlighting the urgent need for better fire retardants. Current fire suppression strategies rely heavily on water, chemical fire retardants, and water-enhancing gels, which use superabsorbent polymers to retain water and adhere to substrates, offering extended fire protection compared to water alone. However, their effectiveness is limited by evaporation and degradation under extreme heat and wind conditions. This study investigates the thermal properties, evaporation dynamics, and fire retardancy mechanisms of a novel polymer-particle (PP) hydrogel with aerogel-forming capabilities. Using a combination of retardancy tests, thermal analysis and modeling, rheological characterization, porosity analysis, and electron microscopy, we analyze how the gelation mechanism, polymer network integrity, and physical interactions of cellulose chains with colloidal silica particles influence fire retardancy and intumescent response. The boiling-induced water vapor expansion and bubble nucleation drive the transformation of the hydrogel into a highly porous, large-volume fire-retardant dome upon rapid heat desiccation, enhancing thermal insulation. By evaluating the retardancy window across different evaporation stages under high heat and wind conditions, this study aims to determine the duration, effectiveness, and governing physical mechanisms of this unique retardant system. These findings provide a framework for designing the next generation of fire retardants with optimized thermal stability and extended protection for wildfire mitigation.
@@ 1 Introduction
Water-enhancing gels (WEGs) have emerged as a promising class of sprayable fire retardants. These materials are engineered to improve water adherence and extend water retention on surfaces by incorporating common superabsorbent polymers.3, 4 By forming a hydrated barrier, WEGs enhance fire resistance and suppression effectiveness through delayed evaporation and reduced heat transfer. However, conventional WEGs rapidly dehydrate under high-temperature and high-wind conditions common during extreme wildfires, significantly curtailing their protective duration in these environments. Studies on the effects of weathering on dehydration and fire performance have underscored the transient nature of these gels, emphasizing the need for formulations that can maintain hydration for extended periods in high-risk scenarios.5 Although the hydration of hydrogels functions as a protective heat sink via endothermic evaporation, the resultant fire-retardant effectiveness is inherently short-lived due to inevitable dehydration.
@@
Recent investigations into intumescent polymers for fire-protective coatings have inspired a new class of hydrogels, which harness polymer-based fire suppression mechanisms to counter climate change-driven wildfire intensification. Intumescent hydrogels represent a significant advancement over conventional intumescent chars, offering improved performance through the formation of protective barrier layers.6, 7, 8, 9, 10 Traditional intumescent coatings typically incorporate ammonium polyphosphate, pentaerythritol, and melamine, which react upon heat exposure to form a char layer that insulates the underlying substrate.11, 12, 13 However, these chars often experience significant shrinkage during combustion, leading to cracking and structural instability that compromise their fire resistance.14, 15 Studies indicate that such chars can shrink by more than 50% during their secondary development, resulting in crack formation and a diminished insulating capability.16 Moreover, while hydrogels offer promise for infrastructure protection, they are typically composed of organic components that inherently exhibit low thermal stability; once desiccated, conventional WEGs lose their effectiveness, limiting their long-term fire protection abilities.
@@
To overcome these limitations, a novel class of water-enhancing hydrogels has been developed that incorporates a hydrogel-to-aerogel transition mechanism to improve thermal stability and flame retardancy.17 This innovative approach leverages rapid intumescence and inorganic particle densification to transform the hydrogel into a solid aerogel dome, thereby providing enhanced mechanical strength and thermal insulation. These polymer-particle (PP) hydrogels are formed through dynamic multivalent interactions between cellulosic biopolymers and colloidal silica particles (CSP), which promote self-assembly into a viscoelastic hydrogel with robust substrate adherence, shear-thinning properties, and extended fire protection.18, 19, 17 Unlike conventional WEGs, which lose efficacy upon desiccation, the heat-activated aerogel transition in PP hydrogels creates a highly porous, thermally insulating silica coating that effectively prevents substrate ignition. This transformation occurs as rapid water evaporation induces silica particle densification and aerogel formation, thereby maintaining a physical thermal barrier under prolonged flame exposure. The PP hydrogel system not only delays substrate charring but also demonstrates superior fire resistance compared to commercial alternatives, introducing a new paradigm in fire science by offering a longer-lasting, high-performance, and environmentally friendly fire-retardant solution.
@@
Building on our previous research reporting the hydrogel-to-aerogel transitions in PP systems,17 we sought to investigate the physical mechanisms underlying PP water-enhancing hydrogel fire protection, analyze its protection window through water loss, and explore the intumescence-driven retardancy at polymer-particle interfaces. The findings provide critical insights into expanding the flame retardancy window of intumescent hydrogels, thereby paving the way for the development of next-generation materials capable of more effectively mitigating catastrophic wildfires.
@@ 2 Results and discussion
In this study we investigate the fire-retardancy window of polymer-particle (PP) hydrogels, focusing on their evaporation dynamics, fire resistance, and foaming behavior. The intumescence effects are examined through the material’s foaming response and microstructural evolution during combustion. To assess performance, we compared a representative PP hydrogel formulation – methyl hydroxyethyl cellulose crosslinked by colloidal silica particles (MHEC/CSP 1-5) – against the commercial water-enhancing gel Phos-Chek® AquaGel-K® (AquaGK). Unlike AquaGK, which comprises of discrete superabsorbent polymer beads, the PP hydrogel forms a continuous, crosslinked dynamic network that exhibits self-healing behavior (Fig. 1(a)) 18, 20, 17. This fundamental difference in microstructure is anticipated to affect both sprayability and fire protection performance.
The selection of MHEC/CSP 1-5 is based on its well-characterized gelation stoichiometry, where colloidal silica particles (LUDOX® TM-50) dynamically interact with methyl hydroxyethyl cellulose (MHEC) chains through adsorption (Figure 1(b)) 21, 18. Improved over earlier formulations for industrial manufacturing 18, this hydrogel exhibits exceptional solution stability due to the minimal conversion of dynamic crosslinks into irreversible covalent bonds, thereby preventing gel aging 17, 20. Previous studies have confirmed the gelation mechanism by tracking changes in storage modulus and loss modulus upon the addition of CSPs to MHEC solutions, validating the formation of a robust, biomimetic, and highly scalable hydrogel network that self-assembles through entropically driven polymer-particle interactions. Furthermore, the commercial availability of both MHEC and CSP ensures practical implementation.
A key advantage of the PP hydrogel over conventional water-enhancing gels, such as AquaGK, is its self-healing property, which maintains structural integrity across the network while preserving sprayability. Rheological analysis via dynamic flow sweeps confirms its shear-thinning behavior, a critical attribute for effective spray application; (Fig. S1) the sprayable nature of PP hydrogels has been verified in an earlier work.
@@ 2.1 Mechanical and material factors governing fire retardancy windows
Increasing the hydration and altering the thermo-physical property of water-based fire-retardant polymers and gels have been recognized as key factors in improving fire retardancy.26 However, for intumescence-based WEGs, such as the MHEC/CSP aerogels reported by Dong et al., the precise mechanisms governing their fire-retardant behavior, retardancy window under evaporation, and intumescence dynamics remain largely unexplored. To address this gap, we systematically examined the behavior of both PP hydrogel and AquaGK systems by preparing formulations at different stages of evaporation (vide infra). This experimental approach simulates the progressive water loss over time, allowing us to assess how evaporation influences both fire retardancy and foaming properties.
@@ 2.1.1 Evaporation Behavior and its Role in Fire Retardancy
Both the MHEC/CSP 1-5 and AquaGK hydrogels exhibited slightly faster evaporation rates than pure water (Fig. 2(b)), suggesting that water retention alone is not the primary determinant of fire retardancy when phenomena such as intumescence, phase transitions, and physical barrier formation are involved. This observation is consistent with previous studies that indicate the thermophysical properties of water droplets and various fire-extinguishing agents – including density, heat capacity, and contact angles – are very similar, resulting in comparable evaporation behaviors 28. The minimal differences in evaporation rates between the two hydrogels imply that conventional WEGs, despite being engineered to extend fire protection via enhanced moisture retention, may ultimately dehydrate similarly to water under high-temperature wildfire conditions. Notably, even though the MHEC/CSP 1-5 hydrogel exhibited a higher rate of water loss than AquaGK at elevated temperatures, it still demonstrated superior flame resistance (vide infra). This reinforces the concept that fire retardancy in PP hydrogels is governed primarily by their heat-activated transition into a thermally insulating aerogel rather than by prolonged water retention alone.
@@ 2.1.2 Gelation Mechanism and Structural Differences
The gelation mechanisms of MHEC/CSP 1-5 and AquaGK are fundamentally different, influencing their behavior upon desiccation and heat exposure. Upon drying, the films formed by AquaGK (Fig. 2(c)(i)) retain a discrete film pattern, consisting of domains of individually crosslinked superabsorbent polymer beads that remain loosely associated within the dried matrix. In contrast, MHEC/CSP 1-5 hydrogels form a cohesive, uniform polymer network, where colloidal silica particles interconnect cellulose chains, maintaining structural integrity and preventing fragmentation (Fig. 2(c)(ii)). This distinction is crucial, as the continuous network structure of MHEC/CSP 1-5 enables stronger substrate adherence and facilitates its mechanically robust aerogel transformation upon heat activation. This is further demonstrated with the combustion process of AquaGK-TMA in Video S2, which is not different from AquaGK, Video S3.
Following heat exposure, MHEC/CSP 1-5 undergoes a striking transition into a porous, thermally insulative aerogel (Fig. 2(c)(iii), top view; Fig. 2(c)(iv), cross-section), exhibiting a distinct foamed morphology compared to the thin, non-intumescent film observed in AquaGK (Fig. 2(c)(i)). This transformation is driven by rapid intumescence and densification of silica particles (Fig. 1(c)(iii)), enabling the aerogel to retain its structural integrity without collapsing, unlike conventional intumescent chars.
@@ 2.1.3 Role of Colloidal Silica in Hydrogel-to-Aerogel Transition
To further investigate the role of colloidal silica particles (CSPs) in aerogel formation, we introduced the same 1-5 concentration ratio of alumina-coated silicon dioxide (LUDOX® TMA) into AquaGK, aiming to induce electrostatic interactions between the negatively charged AquaGK polymeric beads and the positively charged alumina-coated CSPs. This modification was designed to mimic the inorganic-organic interactions found in PP hydrogels and evaluate whether incorporating CSPs into a conventional WEG could replicate the hydrogel-to-aerogel transition observed in MHEC/CSP 1-5.
Despite the addition of CSPs, AquaGK-TMA formulations failed to exhibit the intumescent foaming behavior characteristic of PP hydrogels (Fig. 2(d), Video S1, S2). Instead, the material retained a non-foamed, fragmented residue, similar to unmodified AquaGK. This outcome suggests that the mere presence of CSPs is insufficient to drive the hydrogel-to-aerogel transition; rather, the self-assembled polymer-particle network in MHEC/CSP 1-5 plays a critical role in enabling this transformation. The distinct post-combustion behavior underscores a key limitation of AquaGK’s discretely distributed polymer beads, which lack the network integrity and self-healing properties necessary to support structural expansion during heat activation.
2.1.4 Hydrogel Network Integrity and Intumescence
@@
Furthermore, while surface adhesion contributes to prolonged protection, our findings indicate that additional thermal mechanisms exert a greater influence on overall fire resistance. Specifically, the material’s ability to modify its thermal conductivity and to form secondary insulating structures – such as aerogels and silica shells – upon heat exposure is crucial for enhanced fire protection 17, 29, 8, 9, 10. These structural transformations yield porous, low-conductivity barriers that effectively limit heat transfer, significantly improving insulation and extending the material’s protective lifespan. In contrast, gels that lack these adaptive features depend primarily on evaporative cooling, which is transient and insufficient for long-term fire resistance.
Thus, optimizing fire-retardant gels to promote structural changes upon heating appears to be a more effective approach than relying solely on water retention. Achieving this objective requires a deeper understanding of the underlying mechanisms governing thermally activated aerogel formation and the associated heat transfer properties (vide infra).
2.2 Thermal behavior, intumescence, and retardant mechanism
Thermal conductivity is a directly measurable property of an insulating material and is essential for evaluating the retardancy effectiveness of aerogels. To understand the thermal mechanism behind heat-activated aerogel formation and to calculate the effective thermal conductivity of the aerogels, we model the burn process using the scheme shown in Fig. 3(c).
In this model, the total incident heat from the flame torch, , is balanced by losses due to radiation, convection, and conduction, such that . The input heat is expressed as , where the heat flux from the torch flame is incident on a hydrogel layer of area . The respective losses are defined as . Combining these terms yields the total energy balance as
@@ 2.2 Thermal behavior, intumescence, and retardant mechanism
2.3 Intumescent hydrogel and aerogel morphology
@@
The microstructural evolution of inorganic CSPs within the aerogel system has been previously examined using scanning electron microscopy (SEM).17 However, the role of intumescence in micropore formation and overall microstructural evolution remains unexplored. To address this, additional SEM analyses were performed on cross-sections of aerogel films at different stages of evaporation. The results revealed that heat activation, regardless of the evaporation stage, consistently produced aerogel films composed of multiple layers of CSP with a uniform thickness of 5-10 m (Fig. 4(a)). A notable structural difference emerged when the hydrogel was completely dehydrated into a film without heat activation, as shown in the rightmost image of Fig. 4(a). This dried film exhibited a thickness exceeding 50 m, which is 5-10 times thicker than the heat-activated aerogel layers. This observation elicits that rapid exposure to high temperatures is necessary to achieve the characteristic aerogel microstructure, as opposed to passive drying, which results in a bulkier, non-uniform structure.
@@
Despite these microstructural differences, the macroscopic foaming behavior remained largely unaffected by evaporation. The foaming index, defined as the ratio of post-burn aerogel thickness to pre-burn hydrogel thickness, showed no significant decline even after 75% of the water had evaporated (Fig. 4(b)). This finding contradicts our initial hypothesis that swelling was integral to the intumescence process, instead suggesting that aerogel formation is not hindered by dehydration. Further analysis using Brunauer-Emmett-Teller (BET) surface area measurements confirmed that pore volume and surface area remained consistent across different evaporation stages, with only slight variations. The average pore size remained at 3.4 nm, reinforcing the robustness of the PP hydrogel system in maintaining its flame-retardant properties through substantial water loss that is expected to occur as water evaporates from these materials following application. This pore size classifies the aerogel as a heterogeneous, layered mesoporous structure.30
@@
These findings are particularly significant given that the PP hydrogel system does not rely on conventional intumescent mechanisms, such as acid catalysts or gas-blowing agents, yet still achieves effective fire protection. The ability to maintain structural integrity and thermal insulation even after 75% water loss highlights the resilience and practicality of this hydrogel-based fire-retardant material. The implications of this novel intumescence mechanism, which functions independently of acid and gas-blower chemistry, will be further explored in the next section.

### v2

# Hydrogel-to-Aerogel Transitions in Polymer-Particle Hydrogels Expand the Wildfire Defense Window

@@ Abstract
Keywords: Silica aerogel, dynamic hydrogel, fire retardant, mesoporous, thermal activation
@@ 1 Introduction
Water-enhancing gels (WEGs) have emerged as a promising class of sprayable fire retardants. These materials are engineered to improve water adherence and extend water retention on surfaces by incorporating common superabsorbent polymers.3, 4 By forming a hydrated barrier, WEGs enhance fire resistance and suppression effectiveness through delayed evaporation and reduced heat transfer. Unfortunately, conventional WEGs rapidly dehydrate under high-temperature and high-wind conditions common during extreme wildfires, significantly curtailing their protective duration in these environments on account of their water content-dependent retardant efficacy. Studies on the effects of weathering on dehydration and fire performance have underscored the transient nature of these gels, emphasizing the need for formulations that can maintain performance for extended periods in high-risk scenarios.5 Although the hydration of WEGs functions as a protective heat sink via endothermic evaporation, the resultant fire-retardant effectiveness is inherently short-lived due to inevitable dehydration. Moreover, we believe additional improvements are required to fundamentally change the physical structure and mechanical properties of WEGs, such as surface adherence and spreading properties post application, in addition to water retention, for increasing retardant effectiveness.
@@
Recent investigations into intumescent polymers for fire-protective coatings have inspired a new class of hydrogels, which harness polymer-based fire suppression mechanisms to counter climate change-driven wildfire intensification. Intumescent and aerogel-forming hydrogels represent a significant advancement over conventional intumescent chars, offering improved performance through the formation of protective barrier layers.6, 7, 8, 9, 10 Traditional intumescent coatings include epoxy, acrylic, polyvinyl acetate-based paints, which typically incorporate ammonium polyphosphate, pentaerythritol, and melamine, which react upon heat exposure to form a char layer that insulates the underlying substrate.11, 12, 13 However, these chars often experience significant shrinkage during combustion, leading to cracking and structural instability that compromise their fire resistance.14, 15 Studies indicate that such chars can shrink by more than 50% during their secondary development, resulting in crack formation and a diminished insulating capability.16 Moreover, while hydrogels offer promise for infrastructure protection, they are typically composed of organic components that inherently exhibit low thermal stability; once desiccated, conventional WEGs lose their effectiveness, limiting their long-term fire protection abilities.
@@
To overcome these limitations, a novel class of water-enhancing hydrogels has been developed that incorporates a hydrogel-to-aerogel transition mechanism to improve thermal stability and flame retardancy.17 This innovative approach leverages rapid aerogel formation and inorganic particle densification to transform the hydrogel into a solid aerogel dome, thereby providing enhanced mechanical strength and thermal insulation. These polymer-particle (PP) hydrogels are formed through dynamic multivalent interactions between cellulosic biopolymers and colloidal silica particles (CSP), which promote self-assembly into a viscoelastic hydrogel with robust substrate adherence, shear-thinning properties, and extended fire protection.18, 19, 17 Unlike conventional WEGs, which lose efficacy upon desiccation, the heat-activated aerogel transition in PP hydrogels creates a highly porous, thermally insulating silica coating that effectively prevents substrate ignition. This transformation occurs as rapid water evaporation induces silica particle densification and aerogel formation, thereby maintaining a physical thermal barrier under prolonged flame exposure. The formation of an intumescent aerogel coating and improved mechanical properties, in addition to relaxing the dependence on water content of conventional WEGs, allows the PP hydrogel system to not only delay substrate charring but also demonstrate superior fire resistance compared to commercial alternatives. These benefits introduce a new paradigm in fire science by offering a longer-lasting, high-performance, and environmentally friendly fire-retardant solution.
@@
Building on our previous research reporting the hydrogel-to-aerogel transitions in PP systems,17 we sought to investigate the physical mechanisms underlying PP water-enhancing hydrogel fire protection, analyze its protection window through water loss, and explore the aerogel formation-driven retardancy at polymer-particle interfaces. The findings provide critical insights into expanding the flame retardancy window of intumescent hydrogels, thereby paving the way for the development of next-generation materials capable of more effectively mitigating catastrophic wildfires.
@@ 2 Results and Discussion
In this study we investigate the fire-retardancy window of polymer-particle (PP) hydrogels, focusing on their evaporation dynamics, fire resistance, and foaming behavior. The aerogel effects are examined through the material’s foaming response and microstructural evolution during combustion. To assess performance, we compared a representative PP hydrogel formulation – methyl hydroxyethyl cellulose crosslinked by colloidal silica particles (MHEC/CSP 1-5) – against the commercial water-enhancing gel Phos-Chek® AquaGel-K® (AquaGK). Unlike AquaGK, which comprises of discrete superabsorbent polymer beads, the PP hydrogel forms a continuous, crosslinked dynamic network that exhibits self-healing behavior (Fig. 1(a)) 18, 20, 17. This fundamental difference in microstructure is anticipated to affect both sprayability and fire protection performance.
The selection of MHEC/CSP 1-5 is based on its well-characterized gelation stoichiometry, where colloidal silica particles (LUDOX® TM-50) dynamically interact with methyl hydroxyethyl cellulose (MHEC) chains through adsorption (Figure 1(b)) 21, 18. Improved over earlier formulations for industrial manufacturing 18, this hydrogel exhibits exceptional solution stability due to the minimal conversion of dynamic crosslinks into irreversible covalent bonds, thereby preventing gel aging 17, 20. Previous studies have confirmed the gelation mechanism by tracking changes in storage modulus and loss modulus upon the addition of CSPs to MHEC solutions, validating the formation of a robust, biomimetic, and highly scalable hydrogel network that self-assembles through entropically driven polymer-particle interactions 17, 20. Furthermore, the commercial availability of both MHEC and CSP ensures practical implementation.
@@ 2.1 Mechanical and material factors governing fire retardancy windows
Increasing the hydration and altering the thermo-physical property of water-based fire-retardant polymers and gels have been recognized as key factors in improving fire retardancy.26 However, for aerogel formation-based WEGs, such as the MHEC/CSP aerogels reported by Dong et al., the precise mechanisms governing their fire-retardant behavior, retardancy window under evaporation, and intumescence dynamics remain largely unexplored. To address this gap, we systematically examined the behavior of both PP hydrogel and AquaGK systems by preparing formulations at different stages of evaporation (vide infra). This experimental approach simulates the progressive water loss over time, allowing us to assess how evaporation influences both fire retardancy and foaming properties.
@@ 2.1.1 Evaporation behavior and its role in fire retardancy
Both the MHEC/CSP 1-5 and AquaGK hydrogels exhibited slightly faster evaporation rates than pure water (Fig. 2(b)), suggesting that water retention alone is not the primary determinant of fire retardancy when phenomena such as intumescence, aerogel formation, phase transitions, and physical barrier formation are involved. This observation is consistent with previous studies that indicate the thermophysical properties of water droplets and various fire-extinguishing agents – including density, heat capacity, and contact angles – are very similar, resulting in comparable evaporation behaviors 28. The minimal differences in evaporation rates between the two hydrogels imply that conventional WEGs, despite being engineered to extend fire protection via enhanced moisture retention, may ultimately dehydrate similarly to water under high-temperature wildfire conditions. Notably, even though the MHEC/CSP 1-5 hydrogel exhibited a higher rate of water loss than AquaGK at elevated temperatures, it still demonstrated superior flame resistance (vide infra). This reinforces the concept that fire retardancy in PP hydrogels is governed primarily by their heat-activated transition into a thermally insulating aerogel rather than by prolonged water retention alone.
@@ 2.1.2 Gelation mechanism and structural differences
The gelation mechanisms of MHEC/CSP 1-5 and AquaGK are fundamentally different, influencing their behavior upon desiccation and heat exposure. Upon drying, the films formed by AquaGK (Fig. 2(c)(i)) retain a discrete film pattern, consisting of domains of individually crosslinked superabsorbent polymer beads that remain loosely associated within the dried matrix. In contrast, MHEC/CSP 1-5 hydrogels form a cohesive, uniform polymer network, where colloidal silica particles interconnect cellulose chains, maintaining structural integrity and preventing fragmentation (Fig. 2(c)(ii)). This distinction is crucial, as the continuous network structure of MHEC/CSP 1-5 enables stronger substrate adherence and facilitates its mechanically robust aerogel transformation upon heat activation. This is further demonstrated with the combustion process of AquaGK-TMA in Video S2, which is not different from AquaGK, Video S3. Following heat exposure, MHEC/CSP 1-5 undergoes a striking transition into a porous, thermally insulative aerogel (Fig. 2(c)(iii), top view; Fig. 2(c)(iv), cross-section), exhibiting a distinct foamed morphology compared to the thin, non-foaming film observed in AquaGK (Fig. 2(c)(i)). This transformation is driven by rapid foaming aerogel formation and densification of silica particles (Fig. 1(c)(iii)), enabling the aerogel to retain its structural integrity without collapsing, unlike conventional intumescent chars.
@@ 2.1.3 Role of colloidal silica in hydrogel-to-aerogel transition
To further investigate the role of colloidal silica particles (CSPs) in the observed hydrogel-to-aerogel transition, we introduced the same 1-5 concentration ratio of alumina-coated silicon dioxide (LUDOX® TMA) into AquaGK, aiming to induce electrostatic interactions between the negatively charged AquaGK polymeric beads and the positively charged alumina-coated CSPs. This modification was designed to mimic the inorganic-organic interactions found in PP hydrogels and evaluate whether incorporating CSPs into a conventional WEG material could replicate the hydrogel-to-aerogel transition observed in MHEC/CSP 1-5 hydrogels.
Despite the addition of CSPs, AquaGK-TMA formulations failed to exhibit the intumescent foaming behavior characteristic of PP hydrogels (Fig. 2(d), Video S1, S2). Instead, the material retained a non-foamed, fragmented residue, similar to unmodified AquaGK. This outcome suggests that the mere presence of CSPs is insufficient to drive the hydrogel-to-aerogel transition; rather, the self-assembled polymer-particle network in MHEC/CSP 1-5 plays a critical role in enabling this transformation. The distinct post-combustion behavior underscores a key limitation of AquaGK’s discrete hydrogel beads, which lack the network integrity and self-healing properties necessary to support structural expansion during heat activation. Other possible reasons for this behavior could be the limited foaming of the hydrogels due to slow vapor generation, compounded by thermal effects of the aluminum ions, or the fact ther the superabsorbent polymers used in AquaGK don’t exhibit surfactant-like behaviors.
2.1.4 Hydrogel network integrity and aerogel formation
@@
Furthermore, while surface adhesion contributes to prolonged protection, our findings indicate that additional thermal mechanisms exert a greater influence on overall fire resistance. Specifically, the material’s ability to modify its thermal conductivity and to form secondary insulating structures – such as aerogels and silica shells – upon heat exposure is crucial for enhanced fire protection 17, 29, 8, 9, 10. These structural transformations yield porous, low-conductivity barriers that effectively limit heat transfer, significantly improving insulation and extending the material’s protective lifespan. In contrast, gels that lack these adaptive features depend primarily on evaporative cooling, which is transient and insufficient for long-term fire resistance. Thus, optimizing fire-retardant gels to promote structural changes upon heating appears to be a more effective approach than relying solely on water retention. Achieving this objective requires a deeper understanding of the underlying mechanisms governing thermally activated aerogel formation and the associated heat transfer properties (vide infra).
2.2 Thermal behavior, aerogel formation, and retardant mechanism
Thermal conductivity is a directly measurable property of an insulating material and is essential for evaluating the retardancy effectiveness of aerogels. To understand the thermal mechanism behind heat-activated aerogel formation and to calculate the effective thermal conductivity of the aerogels, we model the burn process using the scheme shown in Fig. 3(c). In this model, the total incident heat from the flame torch, , is balanced by losses due to radiation, convection, and conduction, such that . The input heat is expressed as , where the heat flux from the torch flame is incident on a hydrogel layer of area . The respective losses are defined as . Combining these terms yields the total energy balance as
@@ 2.2 Thermal behavior, aerogel formation, and retardant mechanism
2.3 Hydrogel and aerogel morphology
@@
MHEC/CSP and AquaGK gels exhibit distinct morphological characteristics and flow behaviors (Video S4). AquaGK appears as discrete, bead-like hydrogel particles, whereas MHEC/CSP forms a continuous, cohesive hydrogel network. Beyond these visible differences, the MHEC/CSP hydrogel exhibits smoother flow behavior and superior substrate adherence, both of which are crucial for ease of application and enhanced spreadability for surface-coating applications. The microstructural evolution of inorganic colloidal silica particles (CSPs) within the aerogel system has been previously examined using Scanning Electron Microscopy (SEM), Thermogravimetric Analysis (TGA), Fourier Transform Infrared Spectroscopy (FT-IR), and X-ray Photoelectron Spectroscopy (XPS).17 However, the role of intumescence-led aerification in micropore formation and overall microstructural evolution remains unexplored. To address this, additional SEM analyses were performed on cross-sections of aerogel films at different stages of evaporation. The results revealed that heat activation, regardless of the evaporation stage, consistently produced aerogel films composed of multiple layers of CSP with a uniform thickness of 5-10 m (Fig. 4(a)). A notable structural difference emerged when the hydrogel was completely dehydrated into a film without heat activation, as shown in the rightmost image of Fig. 4(a). This dried film exhibited a thickness exceeding 50 m, which is 5-10 times thicker than the heat-activated aerogel layers. This observation elicits that rapid exposure to high temperatures is necessary to achieve the characteristic aerogel microstructure, as opposed to passive drying, which results in a bulkier, non-uniform structure.
@@
Despite these microstructural differences, the macroscopic foaming behavior remained largely unaffected by evaporation. The foaming index, defined as the ratio of post-burn aerogel thickness to pre-burn hydrogel thickness, showed no significant decline even after 75% of the water had evaporated (Fig. 4(b)). This finding contradicts our initial hypothesis that swelling was integral to the aerogel formation process, instead suggesting that aerogel formation is not hindered by dehydration. Further analysis using Brunauer-Emmett-Teller (BET) surface area measurements confirmed that pore volume and surface area remained consistent across different evaporation stages, with only slight variations. The average pore size remained at 3.4 nm, reinforcing the robustness of the PP hydrogel system in maintaining its flame-retardant properties through substantial water loss that is expected to occur as water evaporates from these materials following application. This pore size classifies the aerogel as a heterogeneous, layered mesoporous structure.30
@@
These findings are particularly significant given that the PP hydrogel system does not rely on conventional intumescent mechanisms, such as acid catalysts or gas-blowing agents, yet still achieves effective fire protection. The ability to maintain structural integrity and thermal insulation even after 75% water loss highlights the resilience and practicality of this hydrogel-based fire-retardant material. The implications of this novel intumescent aerogel formation mechanism, which functions independently of acid and gas-blower chemistry, will be further explored in the next section.